OpenStudy (anonymous):

The length of a rectangle is three times its width. If the perimeter of the rectangle is 48m , find its area.

OpenStudy (anonymous):

2l + 2w = p = 48 l=3w thus w = 6 l = 18 l * w = 108

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

given, l=3b since perimeter = 2(l+b), p=2(3b+b) 48=8b b=6 and l=18

OpenStudy (anonymous):

If L and w is the length and width respectively, then:\[\{L=3w,2L+2w=48\}\] Solve the above for L and w.